OpenStudy (anonymous):

plz help asap very urgent : Find dy/dx , given y =integral( top is exponential^x , bottom is 1 ) (1+surt x )^5 dx

ganeshie8 (ganeshie8):

\[ \dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{x})^5 dx = (1 + \sqrt{e^x})^5 (e^x)'\]

OpenStudy (anonymous):

yes correct explain plz

ganeshie8 (ganeshie8):

use FTC + chainrule

OpenStudy (anonymous):

sorry , wat is ftc

OpenStudy (anonymous):

hello ? explain plz....

ganeshie8 (ganeshie8):

derivative of an integral gives you the function back

ganeshie8 (ganeshie8):

\(\dfrac{d}{dx} \int f = f\)

OpenStudy (anonymous):

ok , so that means there is no long calculation to do , just sub in the top ?

ganeshie8 (ganeshie8):

Yep !

OpenStudy (anonymous):

but why is there another ((ex) at the back ?

ganeshie8 (ganeshie8):

say \(F(x) = \int (1+\sqrt{x})^5 ~dx\)

OpenStudy (anonymous):

ok, continue

ganeshie8 (ganeshie8):

the, \( \int \limits_a^t (1+\sqrt{x})^5 ~dx = F(t) - F(a)\)

ganeshie8 (ganeshie8):

now take the derivative

OpenStudy (anonymous):

oyeah cahin rule u just mentioned , I forgot , haha , thx a thousand :D , actually Im a small mentor in school , then one student came and ask this , now I can help him back :D thanks

ganeshie8 (ganeshie8):

\( \dfrac{d}{dt}\int \limits_a^t (1+\sqrt{x})^5 ~dx = \dfrac{d}{dt} \left(F(t) - F(a)\right) = F'(t) - 0 = F'(t)\)

ganeshie8 (ganeshie8):

oh ok :) since "t" itself is another function, we need to use chain rule

OpenStudy (anonymous):

ya thx I understood :D

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

ftc = fundamental theorem of calculus

ganeshie8 (ganeshie8):

^^ and btw the `variable in d/dx` and the` upper bound` need to match " : \[\dfrac{d}{d \color{red}{x}} \int \limits_a^{\color{red}{x}} f(t) ~ dt = f(x)\]

ganeshie8 (ganeshie8):

if you have a function instead of just "x" : \[\dfrac{d}{d \color{red}{x}} \int \limits_a^{\color{red}{g(x)}} f(t) ~ dt = f(\color{red}{g(x)}) (\color{red}{g(x)} )'\]

OpenStudy (anonymous):

u mean it must be x if its dydx or j if its dydj ?

ganeshie8 (ganeshie8):

exactly ! it has to match, other wise you wud get 0 : d/dx (f(t)) = 0

OpenStudy (anonymous):

huh interesting. Thank you!!! this clears up so much in my multi-var class right now

OpenStudy (anonymous):

lol :P

ganeshie8 (ganeshie8):

:) to make it clear : \[\dfrac{d}{d \color{red}{x}} \int \limits_a^{\color{red}{q}} f(t) ~ dt = 0 \]

OpenStudy (anonymous):

ok im stuck help

OpenStudy (anonymous):

ganeshie u here ??

ganeshie8 (ganeshie8):

im here, is that attachment a different problem ?

OpenStudy (anonymous):

no its the same , thats my calculation and im stuck

ganeshie8 (ganeshie8):

we already finished hte original problem, right ?

OpenStudy (anonymous):

its exactly that question I ask , the attachment is my work , I follow wat u say and do now Im stuck

ganeshie8 (ganeshie8):

\[\dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{x})^5 dx = (1 + \sqrt{e^x})^5 (e^x)' = e^x(1 + \sqrt{e^x})^5 \] we're done.

OpenStudy (anonymous):

u mean I did wrongly ?? did u check my attachment ?

ganeshie8 (ganeshie8):

no, i mean we're done wid the problem. we dont need to do any integration... cuz the integral and derivative eat eachother out by fundamental theorem of calculus

OpenStudy (anonymous):

u chain rule this ? --->> (1+ex−−√)5

ganeshie8 (ganeshie8):

\(e^x(1+\sqrt{e^x})^5\) is the final answer

OpenStudy (anonymous):

I know thats the ans

ganeshie8 (ganeshie8):

hmm

OpenStudy (anonymous):

sorry im still a bit confuse about the extra e^x, u say chain rule , but with wat ?

ganeshie8 (ganeshie8):

chain rule for the top bound

ganeshie8 (ganeshie8):

lets go thru it step by step again

OpenStudy (anonymous):

do u have a paper can u do the chain rule clearly ??

ganeshie8 (ganeshie8):

you want to work below : \[ \dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{x})^5 dx \]

ganeshie8 (ganeshie8):

to make it less confusing, lets change the dummy variable under integration : \[ \dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{t})^5 dt \]

ganeshie8 (ganeshie8):

fine wid this ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

good, say \(F(t) = \int (1+ \sqrt{t})^5 dt ~~~~~\color{red}{*}\)

ganeshie8 (ganeshie8):

\(\implies \) \[ \int \limits _1^{e^x} (1 + \sqrt{t})^5 dt = F(e^x) - F(1) \]

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

still fine ? i have just taken the bounds...

OpenStudy (anonymous):

ya

ganeshie8 (ganeshie8):

next differentiate both sides with respect to x

ganeshie8 (ganeshie8):

\[ \int \limits _1^{e^x} (1 + \sqrt{t})^5 dt = F(e^x) - F(1) \] \(\implies \) \[ \dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{t})^5 dt = \dfrac{d}{dx} \left[F(e^x) - F(1) \right] \]

ganeshie8 (ganeshie8):

still wid me ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

is F (t) this ?

ganeshie8 (ganeshie8):

yes, but we dont need to work any integral here.

OpenStudy (anonymous):

ok ( they cancel eaxh other right ? ) , continue

ganeshie8 (ganeshie8):

integral just disappears by Fundamental thm of calculus, lets do the next step, and it wil be obvious :)

OpenStudy (anonymous):

ok , Im curious and excited :)

ganeshie8 (ganeshie8):

\[\dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{t})^5 dt = \dfrac{d}{dx} \left[F(e^x) - F(1) \right] = \dfrac{d}{dx} \left[F(e^x) \right] - \dfrac{d}{dx} \left[ F(1) \right] \\ = \dfrac{d}{dx} \left[F(e^x) \right] - 0 = \dfrac{d}{dx} \left[F(e^x) \right] \]

ganeshie8 (ganeshie8):

still fine ?

OpenStudy (anonymous):

giv me a moment

OpenStudy (anonymous):

ok, continue

ganeshie8 (ganeshie8):

take ur time :) btw, earlier we defined F(t) as : \(F(t) = \int (1+ \sqrt{t})^5 dt ~~~~~\color{red}{*} \)

ganeshie8 (ganeshie8):

By FTC, taking the derivative gives u the funciton back : \(F(t) = \int (1+ \sqrt{t})^5 dt ~~~~~\color{red}{*} \) \(\implies \dfrac{d}{dt}F(t) = (1+\sqrt{t})^5\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\[ \dfrac{d}{dx}\int \limits _1^{e^x} (1 + \sqrt{t})^5 dt = \dfrac{d}{dx} \left[F(e^x) - F(1) \right] = \dfrac{d}{dx} \left[F(e^x) \right] - \dfrac{d}{dx} \left[ F(1) \right] \\ = \dfrac{d}{dx} \left[F(e^x) \right] - 0 = \dfrac{d}{dx} \left[F(e^x) \right] \]

ganeshie8 (ganeshie8):

use the previous result( \(\color{red}{*}\)) to evaluate above ^

OpenStudy (anonymous):

ok, but there is no second e^x?

ganeshie8 (ganeshie8):

good question, say \(e^x = u\)

ganeshie8 (ganeshie8):

By chain rule : \[ \dfrac{d}{dx} \left[F(e^x) \right] = \dfrac{d}{dx} \left[F(u) \right] * \dfrac{d}{dx} [u] \]

ganeshie8 (ganeshie8):

\[= (1+\sqrt{u})^5 \dfrac{d}{dx} [u]\]

ganeshie8 (ganeshie8):

plugin the value of \(u\)

OpenStudy (anonymous):

a moment plz.... digesting

ganeshie8 (ganeshie8):

k

OpenStudy (anonymous):

if u chain rule , isnt the power 5 suppose to bring to the front and the original minus by 1 ?

ganeshie8 (ganeshie8):

nope, \[F(t) = \int (1+ \sqrt{t})^5 dt ~~~~~\color{red}{*} \] \[ \implies \dfrac{d}{dt}F(t) = (1+\sqrt{t})^5 \]

ganeshie8 (ganeshie8):

^^thats directly by Fundamental theorem of caluclus, you okay wid that right ?

OpenStudy (anonymous):

im not sure about the chan rule , cuz mine is completely different , I let u see , wait I snap photo

ganeshie8 (ganeshie8):

are you fine wid Fundamental theorem of calculus above ?

ganeshie8 (ganeshie8):

lets finish this, before moving to seeing ur work

ganeshie8 (ganeshie8):

too many things = too many confusions lol

OpenStudy (anonymous):

wait , u just have a look at this first

ganeshie8 (ganeshie8):

looks u have attached the wrong pic, check once

OpenStudy (anonymous):

yes its this , look at the chain rule

OpenStudy (anonymous):

no.8

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

why it brings the power to the front ?

ganeshie8 (ganeshie8):

thats the chain rule for taking derivative of a specific function of type : \([f(x) ]^n\)

OpenStudy (anonymous):

ok, which no. should I be looking at ?

OpenStudy (anonymous):

this is not in the table ?

ganeshie8 (ganeshie8):

a more general chain rule : \[y = f(g(x)) \implies \dfrac{dy}{dx} = f'(g(x)) \times g'(x)\]